OpenStudy (anonymous):

what is the answer to 5(2x+1)=3(5x-4)

OpenStudy (anonymous):

\(\Large 5(2x+1)=3(5x-4)\) \(\Large 10x+5=15x-12\) \(\Large 15x-10x=5+12\) \(\Large 5x=17\) \(\Large x=\frac{17}{5}

OpenStudy (anonymous):

\(\Large x=\frac{17}{5}\)